# Quantifying artificial intelligence through algebraic generalization

### v1

@@ Abstract
The rapid development of modern artificial intelligence (AI) systems has created an urgent need for their scientific quantification. While their fluency across a variety of domains is impressive, modern AI systems fall short on tests requiring symbolic processing and abstraction – a glaring limitation given the necessity for interpretable and reliable technology. Despite a surge of reasoning benchmarks emerging from the academic community, no comprehensive and theoretically-motivated framework exists to quantify reasoning (and more generally, symbolic ability) in AI systems. Here, we adopt a framework from computational complexity theory to explicitly quantify symbolic generalization: algebraic circuit complexity. Many symbolic reasoning problems can be recast as algebraic expressions. Thus, algebraic circuit complexity theory – the study of algebraic expressions as circuit models (i.e., directed acyclic graphs) – is a natural framework to study the complexity of symbolic computation. The tools of algebraic circuit complexity enable the study of generalization by defining benchmarks in terms of their complexity-theoretic properties (i.e., the difficulty of a problem). Moreover, algebraic circuits are generic mathematical objects; for a given algebraic circuit, an arbitrarily large number of samples can be generated for a specific circuit, making it an optimal testbed for the data-hungry machine learning algorithms that are used today. Here, we adopt tools from algebraic circuit complexity theory, apply it to formalize a science of symbolic generalization, and address key theoretical and empirical challenges for its successful application to AI science and its impact on the broader community.
@@ Introduction
The ability to reason algebraically is often considered a hallmark of human intelligence Marcus, 2003. The recent evolution of modern artificial intelligence (AI) systems and large language models (LLMs) has led to the speculation that these systems may also reason algebraically Bubeck et al., 2023, Wei et al., 2022a, Webb et al., 2023. Yet due to challenge of evaluating large models trained on massive pretraining datasets Kim et al., 2022, it is difficult to evaluate whether such models are truly exhibiting algebraic reasoning abilities, or whether they instead regurgitate plausible text from their pretraining data Schaeffer et al., 2023, Wu et al., 2024. This ambiguity has led to a deluge of symbolic reasoning benchmarks Lake and Baroni, 2018, Hupkes et al., 2020, Hudson and Manning, 2019, Yang et al., 2018, Ito et al., 2024, Johnson et al., 2017, Clark et al., 2019, Kim and Linzen, 2020, Ruis et al., 2020, Keysers et al., 2020. Despite these efforts, objectively quantifying the complexity of reasoning problems is difficult; most of these experiments are ad hoc, and designed without a framework to quantify complexity. However, approaches in computational complexity theory, a field within theoretical computer science, have made it possible to explicitly measure a problem’s algorithmic difficulty, enabling the design of generalization tests rooted in quantifiable measures of complexity. Here, we adopt a branch of computational complexity theory – algebraic circuit complexity – to provide a parsimonious set of problems and approaches to rigorously study symbolic computation in machine learning.
@@
Recently, there has been increased interest in studying AI models through arithmetic and compositional tasks Kudo et al., 2023, McLeish et al., 2024, Zhou et al., 2024a, Shen et al., 2024, Saxton et al., 2019, Lee et al., 2023, Dziri et al., 2023, McCoy et al., 2023, Zhou et al., 2024b, Ito et al., 2024, Hupkes et al., 2020, Sinha et al., 2024. Compositional tasks are problems that are generated by recombining a basis set of atomic elements to form a variety of task combinations. (Arithmetic problems are compositional; they are composed of atomic elements (numbers and operators), and can be recomposed to generate novel expressions and problems.) These tasks are good reasoning benchmarks because they require 1) abstraction, 2) logical application of rules or axioms, and 3) precise problem-solving and rigor. Critically, these paradigms have provided reliable ways to elicit failure modes in transformer-based AI models for specific forms of symbolic generalization. For example, a number of studies have demonstrated the difficulty of “length generalization” – generalizing to problems of longer length than seen during training Kazemnejad et al., 2023, McLeish et al., 2024, Zhou et al., 2024b, Shen et al., 2024. Hupkes et al. 2020 and Hupkes et al. 2023 also introduced various notions (e.g., systematicity and productivity) in an effort to taxonomize different forms of linguistic generalization. While incredibly useful for linguistics, it is unclear how these concepts generically apply beyond natural language processing. By contrast, the formalisms from algebraic circuit complexity theory provide a set of parsimonious mathematical tools that can be applied to quantify symbolic generalization, and are agnostic to specific domains, such as linguistics. Moreover, algebraic circuit complexity theory provides an encompassing theoretical framework for the increasingly popular, yet nascent empirical evaluations in AI systems that use arithmetic tasks Kudo et al., 2023, McLeish et al., 2024, Deletang et al., 2022, Zhou et al., 2023, Shen et al., 2024, Saxton et al., 2019, Lee et al., 2023, Dziri et al., 2023, Ruoss et al., 2023, Jelassi et al., 2023, Wang et al., 2021, Nogueira et al., 2021.
@@
A large class of symbolic problems can be studied with algebraic expressions Shpilka and Yehudayoff, 2010, Bürgisser et al., 2013. Algebraic circuit complexity theory formalizes algebraic expressions as circuit models (i.e., directed acyclic graphs; Fig. 1). This formalization is well-established in computational complexity theory, the branch of theoretical computer science concerned with quantifying the difficulty of computational problems and the resources required to solve them Arora and Barak, 2009. Importantly, formalizing computational problems in terms of circuits is the leading approach to empirically quantify their complexity. Unlike other notions of complexity, such as Kolmogorov Complexity in algorithmic information theory (which is incomputable), notions developed in complexity theory for circuits are explicitly computable and determined by their shape and structure Wyeth and Sturtivant, 2023. Thus, the tools of algebraic circuit complexity can formalize notions of generalization by defining benchmarks in terms of their circuit properties. Furthermore, algebraic circuits are generic mathematical objects; they can be represented from a variety of mathematical perspectives (geometry, topology, etc.), providing useful interpretations in other domains. Algebraic circuits are therefore well-situated to addressing problems in symbolic machine learning – the problems are computable, large datasets can be straightforwardly generated from circuit specifications, and new models can be developed that address specific failure modes within this framework. In the ensuing sections, we provide a blueprint for the successful adoption of algebraic circuit complexity for machine learning problems; we introduce the core components of algebraic circuits, address how they can be leveraged to study symbolic generalization, and discuss several key open theoretical and empirical challenges.
@@
Algebraic circuit complexity studies algebraic expressions as computable circuit models. There has been a significant amount of recent machine learning research studying arithmetic generalization, a key symbolic ability. While important and insightful, most of those studies primarily focus on a restricted set of algebraic problems, which we will illustrate later Dziri et al., 2023, Kudo et al., 2023, McLeish et al., 2024, Shen et al., 2024, Jelassi et al., 2023. Below, we provide definitions of algebraic circuits that will place prior work within a broader mathematical framework. Our goal is to provide the tools to quantify model generalization through circuit complexity, rather than ad hoc goals such as, for example, “length generalization”.
Definition. (For a more formal definition, see Shpilka and Yehudayoff 2010 or Bürgisser et al. 2013.) An algebraic circuit $C$ represents a polynomial expression as a directed acyclic graph, comprised of gates $v$ and edges $e$. Input gates, $v_{X}$ and $v_{\mathbb{F}}$, are respectively defined as either variables (e.g., $X=\{x_{1},...,x_{n}\}$) or elements in a field $\mathbb{F}$ (e.g., $\mathbb{R})$. Input gates have fan-in (in-degree) of 0. All other gates are operators: a sum gate ($v_{+}$) or a product gate ($v_{\times}$). Here we restrict the fan-in of operator gates to 2, as is standard Shpilka and Yehudayoff, 2010. For our purposes, this ensures that there is the same number of gates in a circuit model with its corresponding representation as a string of tokens, which simplifies downstream analyses (e.g., the analysis in Fig. 6). Operator gates have a fan-out (out-degree) of either 1 or 0. If the fan-out or an operator gate is 0, then it is the output gate of that polynomial.
Properties. An algebraic circuit $C$ has two main complexity measures: size $s$ and depth $d$ (see Table 1 for a summary of all properties). The size $s$ of a circuit refers to the number of edges $e$ in $C$. The depth $d$ of a circuit refers to the longest path from an input gate to the output gate. We can denote a sub-circuit $C_{v}$ of $C$, which computes the polynomial $f_{v}$ rooted at gate $v$. Another important property of an algebraic circuit is its degree (i.e., the degree of a polynomial). The degree of a circuit (or a sub-circuit) can be computed by measuring the degree of the gate $v$, denoted $deg(v)$. Elements in a field $\mathbb{F}$ are of degree 0, input variables $x\in X$ have degree 1, the degree of an sum gate (++) is determined by the degrees of its inputs $u$, $v$ such that $deg(u+v)=max(deg(u),deg(v))$, and the degree of a product gate ($\times$) is determined by $deg(u\cdot v)=deg(u)+deg(v)$.
@@
Figure 1 provides a few simple examples of algebraic circuits. A circuit representation of an algebraic expression provides a concrete algorithm for computing a polynomial, with some circuits requiring more computation than others (e.g., determined by size or depth). Furthermore, a circuit description can provide explicit differences in required computation for polynomials that are mathematically equivalent (e.g., see Fig. 1C,D). Such a formalization can be useful to study how different representations of equivalent algebraic expressions can lead to different levels of generalization. The recent studies in arithmetic generalization in AI models focus on the simplest circuit representations of algebraic problems (problems analogous to Fig. 1A, but varying the magnitude of the field elements, e.g., train on small digit numbers, test on large digit numbers). The language of circuit complexity can provide more interesting and flexible ways to investigate generalization across circuit complexity classes.
@@
A recent study concluded that LLMs’ behavior is a function of the problems they are trained to solve McCoy et al., 2023. It is difficult to identify what the “problems” are when the pretraining corpus is natural language text derived from the internet. A theoretically coherent alternative to quantify AI systems is to select problems where 1) difficulty is quantifiable, and 2) arbitrarily large datasets can be systematically generated. Algebraic circuits satisfy both of these constraints.
Though recent papers have demonstrated an increased interest in using arithmetic tasks to quantify generalization, preliminary approaches have been theoretically limited. As alluded to above, many of the recent papers evaluating “length generalization” in arithmetic tasks train on two variable addition (or multiplication), and test on the same circuit class but sample field elements that are larger in magnitude (Fig. 2A) Zhou et al., 2024a, Shen et al., 2024, Nogueira et al., 2021, McLeish et al., 2024, Zhou et al., 2024b, Lee et al., 2023, Jelassi et al., 2023. Others focus on a more complex form of generalization, i.e., addition or multiplication on problems with a greater number of variables, which is equivalent to increasing the circuit size and depth (Fig. 2B) (e.g., modular arithmetic task; Deletang et al. 2022). While useful and informative, both these tests of generalization scratch the surface of generalization metrics that can be devised with algebraic circuits.
Figure 2: Commonly-used AI evaluations for “length generalization” with arithmetic tasks. A) The predominant form of “length generalization” in transformers is evaluating performance on addition or multiplication problems with larger integers than seen during training Shen et al., 2024, Jelassi et al., 2023, McLeish et al., 2024, Zhou et al., 2024b, Lee et al., 2023, Zhou et al., 2024a. This would be conceptually equivalent to a situation in which the circuit size and depth are fixed, but the sampling of input gates (i.e., field elements) differs across training and testing sets. The notion of length generalization is specific to the context of transformers, given that larger digits require a larger context window. B) Another form of “length generalization” studied in the literature is to generalize to arithmetic problems with more operands (variables) than seen in the training set Deletang et al., 2022. From a circuit complexity perspective, these two approaches are distinct problems.
@@
We first define symbolic generalization (Box 1), and consider the importance of learning algebraically (or compositionally). The ability to symbolically generalize is one of the most challenging problems in AI, and is a requirement for robust reasoning and planning. Like algebraic circuits, several prior papers formulated symbolic tasks as computing a path through a directed graph or circuit Dziri et al., 2023, Yang et al., 2018, Johnson et al., 2017. Importantly, learning a symbolic task requires 1) decomposing and abstracting the individual functions (e.g., nodes) of a circuit, and understanding how they can be recomposed via edges to produce novel circuits (e.g., function composition). Algebra is a natural language to study symbolic generalization as it can be encoded as a circuit, and requires both abstraction and function composition. Moreover, in contrast to other compositional approaches (e.g., regular or context-free grammars; Deletang et al. 2022, Chomsky 1956), algebraic problems encompass an infinite vocabulary (e.g., $\mathbb{R}$) with an expressive grammar determined by its axioms, making it well-suited for machine learning algorithms that require many training samples. In the following section, we propose a framework of generalization in terms of algebraic circuits.
@@
Box 1. A formal sketch of symbolic generalization. Let $B$ be a basis set of elements. Let $U_{1}=B$ and $U_{i+1}=\{u\mid(u_{1},u_{2}\in U_{\leq i})\text{ \& }(u=u_{1}\circ u_{2})\}$, where $\circ$ stands for any computable composition operator on $U_{i}$ (akin to production rules in context-free grammars or operations in a field). Then, we define some universe set $U=\bigcup_{i\in\mathbb{N}}U_{i}$ We define a function (task) $T$ on $U$ such that $T:U\rightarrow R$ (see below for tasks $T$ for algebraic circuits). $T$ is a compositional function, where $\forall u_{1},u_{2}\in U$, $T(u_{1}\circ u_{2})=T(u_{1})\ast T(u_{2})$, and $\ast$ stands for any computable composition on $R$.
@@
Let $D_{train}$ and $D_{test}$ be distributions over $U$, and $supp(D_{train}|B)$ and $supp(D_{test}|B)$ to be the support of the basis set elements in $D_{train}$ and $D_{test}$, respectively. We restrict $D_{test}$ to be a distribution in which $supp(D_{test}|B)\subseteq supp(D_{train}|B)$. Additionally, if a composable function $\circ$ is represented in $D_{test}$ (i.e., $u\in D_{test}$ and $u=b_{1}\circ b_{2}$), then we require that $\circ$ is also represented in $D_{train}$. We include these as requirements for symbolic generalization; it would be challenging, if not impossible, to generalize to samples in $D_{test}$ if not all basis elements and composition operators were provided in $D_{train}$.
@@
We define symbolic/compositional generalization as the evaluation of a learned model $M$ (e.g., a neural network) where $Pr_{x\sim D_{test}}[M(x)=T(x)]>1-\epsilon$ (1) Here, $M$ is only optimized from samples $x\sim D_{train}$, and $\epsilon$ is the evaluation error of $M$ on $x\sim D_{test}$. A strong form of symbolic generalization over algebraic circuits would be to satisfy Equation 1 for all valid partitions of $D_{train}$ and $D_{test}$ (with the requisite support of basis sets and composition functions) in $U$ for uniformly small values of $\epsilon$.
@@
The language of algebraic circuits provides quantitative metrics to formalize generalization. In particular, generalization of AI systems can be quantified in terms of circuit divergence – the divergence of circuit parameters between algebraic circuits employed during training and testing (Fig. 3). In other words, given a model $M$, we want to characterize its generalization performance $M(C_{test}|\mathcal{C})$, where $C_{test}$ is a test circuit, and $\mathcal{C}=\{C_{1},...,C_{k}\}$ is a family of training circuits. We define circuit divergence as the difference between quantifiable circuit properties between train and test distributions. Here, we emphasize five important circuit properties to measure divergences: the sampling of a field $\mathbb{F}$, the number of variables $\lvert X\rvert$, size $s$, depth $d$, and the polynomial degree. By quantifying the properties of circuits in both the training and testing sets, circuit divergence can be explicitly measured along these dimensions. In the following sections, we will provide several examples where tests of generalization can be constructed through the manipulation of circuit divergence.
Figure 3: Symbolic capabilities of modern AI systems and architectures can be studied with algebraic circuits. A) A set of problems can be identified (and sampled) from algebraic circuits. B) Given a family of circuits, we can identify circuit divergences – the divergence of different circuit properties, like size and depth – to design train datasets and test datasets to evaluate the algebraic generalization ability of a model. C,D) We can evaluate a model (or set of models) across these dataset splits, and quantify their performance according to circuit divergences. This AI evaluation lifecycle allows us to iterate and refine hypotheses regarding the degree to which a model can generalize to a class of algebraic problems of particular complexity. (Note that circuit divergences can be measured beyond depth and size.)
@@
In this section, we illustrate the flexibility of algebraic circuits in designing meaningful AI benchmarks. To demonstrate the generality of algebraic circuits, we begin by providing benchmarks that are analogous to popular tests of compositional generalization: systematicity and productivity Hupkes et al., 2020. We subsequently introduce novel, more challenging problems that can be used to evaluate more general types of symbolic computation and abstraction. Finally, we demonstrate an approach to reformulate tests of algebraic generalization for pretrained LLMs.
@@
There has been recent interest in using compositional paradigms to study the symbolic ability of machine learning models Hupkes et al., 2020, Lake and Baroni, 2018, 2023, Ruis and Lake, 2022, Dziri et al., 2023, Ontanon et al., 2022, Csordás et al., 2021, Ito et al., 2024, Klinger et al., 2023, Poggio and Fraser, 2024, Ram et al., 2024. Here we demonstrate direct links to popular forms of compositional generalization using algebraic circuits.

### v2

# Quantifying artificial intelligence through algorithmic generalization

@@ Abstract
The rapid development of artificial intelligence (AI) systems has created an urgent need for their scientific quantification. While their fluency across a variety of domains is impressive, AI systems fall short on tests requiring algorithmic reasoning – a glaring limitation given the necessity for interpretable and reliable technology. Despite a surge of reasoning benchmarks emerging from the academic community, no theoretical framework exists to quantify algorithmic reasoning in AI systems. Here, we adopt a framework from computational complexity theory to quantify algorithmic generalization using algebraic expressions: algebraic circuit complexity. Algebraic circuit complexity theory – the study of algebraic expressions as circuit models – is a natural framework to study the complexity of algorithmic computation. Algebraic circuit complexity enables the study of generalization by defining benchmarks in terms of the computational requirements to solve a problem. Moreover, algebraic circuits are generic mathematical objects; an arbitrarily large number of samples can be generated for a specified circuit, making it an ideal experimental sandbox for the data-hungry models that are used today. In this Perspective, we adopt tools from algebraic circuit complexity, apply them to formalize a science of algorithmic generalization, and address key challenges for its successful application to AI science.
@@ Introduction
The recent evolution of modern artificial intelligence (AI) systems and large language models (LLMs) has led to the speculation that these systems may reason 1, 2, 3, 4. Yet due to challenge of evaluating large models trained on massive pretraining datasets 5, it is difficult to evaluate whether such models are truly exhibiting algorithmic reasoning abilities, or whether they instead regurgitate plausible text from their pretraining data 6, 7. This ambiguity has led to a deluge of reasoning benchmarks 8, 9, 10, 11, 12, 13, 14, 15, 16, 17. Despite these efforts, objectively quantifying the complexity of reasoning problems is difficult; most of these experiments are ad hoc, and designed without a framework to quantify and verify the algorithmic complexity of reasoning problems. However, approaches in computational complexity theory, a field within theoretical computer science, have made it possible to explicitly measure a problem’s algorithmic difficulty, paving the way for generalization tests rooted in quantifiable measures of complexity. In this Perspective, we bridge algorithmic reasoning with a decades-old branch of computational complexity – circuit complexity theory – to provide a theoretical link to studying the complexity of algorithmic computation in modern AI systems.
@@
Recently, there has been increased interest in studying modern AI models through arithmetic and compositional tasks 18, 19, 20, 21, 22, 23, 24, 25, 26, 12, 9, 27. Compositional tasks are problems rooted in a long history from the early 20th century28, 29 that are generated by recombining a basis set of atomic elements to form a variety of task combinations (for a review, see Russin et al. 30). (Arithmetic problems are compositional; they are composed of atomic elements (numbers and operators), and can be recomposed to generate novel expressions and problems.) For modern AI systems, compositional tasks can serve as useful reasoning benchmarks since they require 1) abstraction, 2) logical and verifiable application of rules or axioms, and 3) precise problem-solving and rigor. Critically, these paradigms have provided reliable ways to elicit failure modes in transformer-based AI models for specific forms of compositional generalization. For example, a number of studies have demonstrated the difficulty of “length generalization” – generalizing to problems of longer length than seen during training 31, 19, 26, 21. Other researchers have also introduced various notions (e.g., systematicity and productivity) in an effort to taxonomize different forms of compositionality 9, 32, 33, 34, 35. By contrast, the formalisms from circuit complexity theory provide a set of tools that can be applied to quantify algorithmic generalization – generalization over algorithms specified by circuits and their associated measures of complexity, such as space or time complexity. (Structural properties of circuits correspond directly to algorithmic requirements to compute that problem.) Moreover, formalizing problems through a circuit complexity framework provides a theoretically grounded framework for the increasingly popular, yet nascent empirical evaluations in AI systems that use arithmetic and compositional tasks 18, 19, 36, 37, 21, 22, 23, 24, 38, 39, 40, 41. Though we focus on formalizing algebraic problems through the lens of circuit complexity (i.e., algebraic circuit complexity) due to the widespread use of arithmetic problems to evaluate modern AI systems, the more general framework of circuits can be similarly extended to other algorithmic problems.
@@
A large class of algorithmic problems can be studied with algebraic expressions 42, 43. Algebraic circuit complexity theory formalizes the evaluation of algebraic expressions through algorithms encoded as a computable circuits (i.e., directed acyclic graphs; Fig. 1). This formalization is well-established in computational complexity theory, the branch of theoretical computer science concerned with quantifying the difficulty of computational algorithms and the resources required to solve them 44. Importantly, formalizing computational problems in terms of circuits is the leading approach to empirically quantify their complexity. Unlike other notions of complexity, such as Kolmogorov Complexity in algorithmic information theory (which is incomputable), notions developed in complexity theory for circuits are explicitly computable and determined by their shape and structure 45. Thus, the tools of circuit complexity can formalize notions of generalization over algorithms by defining benchmarks in terms of their circuit properties. Furthermore, algebraic circuits are generic mathematical objects; they can be represented from a variety of mathematical perspectives (geometry, topology, etc.), providing useful interpretations in other domains. Algebraic circuits are therefore well-situated to investigate algorithmic generalization – the problems are computable and verifiable, large datasets can be straightforwardly generated from circuit specifications, and new models can be developed that address specific failure modes within this framework. In the ensuing sections, we provide a blueprint for the successful adoption of algebraic circuit complexity as a gateway into studying algorithmic generalization more broadly. We introduce the core components of algebraic circuits, address how they can be leveraged to study algorithmic generalization, and discuss several key open theoretical and empirical challenges.
@@
Algebraic circuit complexity studies algebraic expressions as computable circuit models. There has been a significant amount of recent machine learning research studying arithmetic generalization, a key algorithmic ability. While important and insightful, most of those studies primarily focus on a restricted set of algebraic problems, which we will illustrate later 24, 18, 19, 21, 39. Below, we provide definitions of algebraic circuits that will place prior work within a broader mathematical framework. Our goal is to provide the tools to quantify model generalization through circuit complexity, rather than ad hoc goals such as, for example, “length generalization”.
Definition. (For a more formal definition, see Shpilka and Yehudayoff 42 or Bürgisser et al. 43.) An algebraic circuit $C$ represents a polynomial expression as a directed acyclic graph, comprised of gates $v$ and edges $e$. Input gates, $v_{X}$ and $v_{\mathbb{F}}$, are respectively defined as either variables (e.g., $X=\{x_{1},...,x_{n}\}$) or elements in a field $\mathbb{F}$ (e.g., $\mathbb{R})$. Input gates have fan-in (in-degree) of 0. All other gates are operators: a sum gate ($v_{+}$) or a product gate ($v_{\times}$). In this Perspective, we restrict the fan-in of operator gates to 2, as is standard 42. For our purposes, this ensures that there is the same number of gates in a circuit model with its corresponding representation as a string of tokens, which simplifies downstream analyses (e.g., the analysis in Fig. 6). Operator gates have a fan-out (out-degree) of either 1 or 0. If the fan-out of an operator gate is 0, then it is the output gate of that polynomial.
Properties. An algebraic circuit $C$ has two main algorithmic complexity measures: size $s$ and depth $d$ (see Table 1 for a summary of all properties). The size $s$ of a circuit refers to the number of edges $e$ in $C$, and corresponds to algorithmic space complexity. The depth $d$ of a circuit refers to the longest path from an input gate to the output gate, and corresponds to algorithmic time complexity. We can denote a sub-circuit $C_{v}$ of $C$, which computes the polynomial $f_{v}$ rooted at gate $v$. Another important property of an algebraic circuit is its degree (i.e., the degree of a polynomial). The degree of a circuit (or a sub-circuit) can be computed by measuring the degree of the gate $v$, denoted $deg(v)$. Elements in a field $\mathbb{F}$ are of degree 0, input variables $x\in X$ have degree 1, the degree of an sum gate ($+$) is determined by the degrees of its inputs $u$, $v$ such that $deg(u+v)=max(deg(u),deg(v))$, and the degree of a product gate ($\times$) is determined by $deg(u\cdot v)=deg(u)+deg(v)$.
@@
Figure 1 provides a few simple examples of algebraic circuits. A circuit representation of an algebraic expression provides a concrete algorithm for computing a polynomial, with some circuits requiring more computation than others (e.g., determined by size or depth). Furthermore, a circuit description can provide explicit differences in required computation for polynomials that are mathematically equivalent (e.g., see Fig. 1C,D). Such a formalization can be useful to study how different representations of equivalent algebraic expressions can lead to different levels of generalization, or how syntax relates to semantics in formal languages. The recent studies in arithmetic generalization in AI models focus on the simplest circuit representations of algebraic problems (problems analogous to Fig. 1A, but varying the magnitude of the field elements, e.g., train on small digit numbers, test on large digit numbers). The language of circuit complexity can provide more interesting and flexible ways to investigate generalization across circuit complexity classes.
@@
A recent study concluded that LLMs’ behavior is a function of the problems they are trained to solve 25. It is difficult to identify what the “problems” are when the pretraining corpus is natural language text derived from the internet. A theoretically coherent alternative to quantify AI systems is to select problems where 1) complexity is quantifiable, and 2) arbitrarily large datasets can be systematically generated. Algebraic circuits satisfy both of these constraints.
Though recent papers have demonstrated an increased interest in using arithmetic tasks to quantify generalization, preliminary approaches have been theoretically limited. As alluded to above, many of the recent papers evaluating “length generalization” in arithmetic tasks train on two variable addition (or multiplication), and test on the same circuit class but sample field elements that are larger in magnitude (Fig. 2A) 20, 21, 41, 19, 26, 23, 39. Others focus on a more complex form of generalization, i.e., addition or multiplication on problems with a greater number of variables, which is equivalent to increasing the circuit size and depth (Fig. 2B) (e.g., modular arithmetic task; Deletang et al. 36). While useful and informative, both these tests of generalization scratch the surface of generalization metrics that can be devised with algebraic circuits.
Figure 2: Commonly-used AI evaluations for “length generalization” with arithmetic tasks. A) The predominant form of “length generalization” in transformers is evaluating performance on addition or multiplication problems with larger integers than seen during training 21, 39, 19, 26, 23, 20. This would be conceptually equivalent to a situation in which the circuit size and depth are fixed, but the sampling of input gates (i.e., field elements) differs across training and testing sets. (However, see discussion below in Open theoretical and empirical challenges on the nuances of computing long addition.) The notion of length generalization is specific to the context of transformers, given that larger digits require a larger context window. B) Another form of “length generalization” studied in the literature is to generalize to arithmetic problems with more operands (variables) than seen in the training set 36. From a circuit complexity perspective, these two approaches are distinct problems.
@@
We first define algorithmic generalization (Box 1), and consider the importance of learning algorithmically. The ability to generalize algorithmically is one of the most challenging problems in AI, and is a requirement for robust reasoning and planning. Like algebraic circuits, several prior papers formulated algorithmic tasks as computing a path through a directed graph or circuit 24, 11, 13, 46. Importantly, learning an algorithmic task requires 1) decomposing and abstracting the individual functions (e.g., nodes) of a circuit, and understanding how they can be recomposed via edges to produce novel circuits (e.g., function composition). Algebra is a natural language to study algorithmic generalization as it can be encoded as a circuit, and requires both abstraction and function composition. Moreover, in contrast to other compositional approaches (e.g., regular or context-free grammars; Deletang et al. 36, Chomsky 47), algebraic problems encompass an infinite vocabulary (e.g., $\mathbb{R}$) with an expressive grammar determined by its axioms, making it well-suited for machine learning algorithms that require many training samples. In the following section, we propose an algorithmic generalization framework in terms of algebraic circuits.
@@
Box 1. A formal sketch of algorithmic generalization. Let $B$ be a basis set of elements. Let $U_{1}=B$ and $U_{i+1}=\{u\mid(u_{1},u_{2}\in U_{\leq i})\text{ \& }(u=u_{1}\circ u_{2})\}$, where $\circ$ stands for any computable composition operator on $U_{i}$ (akin to production rules in context-free grammars or operations in a field). Then, we define some universe set $U=\bigcup_{i\in\mathbb{N}}U_{i}$ We define a function (task) $T$ on $U$ such that $T:U\rightarrow R$ (see below for tasks $T$ for algebraic circuits). $T$ is a compositional function, where $\forall u_{1},u_{2}\in U$, $T(u_{1}\circ u_{2})=T(u_{1})\ast T(u_{2})$, and $\ast$ stands for any computable composition on $R$.
@@
Let $D_{train}$ and $D_{test}$ be distributions over $U$, and $supp(D_{train}|B)$ and $supp(D_{test}|B)$ to be the support of the basis set elements in $D_{train}$ and $D_{test}$, respectively. We restrict $D_{test}$ to be a distribution in which $supp(D_{test}|B)\subseteq supp(D_{train}|B)$. Additionally, if a composable function $\circ$ is represented in $D_{test}$ (i.e., $u\in D_{test}$ and $u=b_{1}\circ b_{2}$), then we require that $\circ$ is also represented in $D_{train}$. We include these as requirements for algorithmic generalization; it would be challenging, if not impossible, to generalize to samples in $D_{test}$ if not all basis elements and composition operators were provided in $D_{train}$.
@@
We define algorithmic generalization as the evaluation of a learned model $M$ (e.g., a neural network) where $Pr_{x\sim D_{test}}[M(x)=T(x)]>1-\epsilon$ (1) Here, $M$ is only optimized from samples $x\sim D_{train}$, and $\epsilon$ is the evaluation error of $M$ on $x\sim D_{test}$. A strong form of algorithmic generalization over algebraic circuits would be to satisfy Equation 1 for all valid partitions of $D_{train}$ and $D_{test}$ (with the requisite support of basis sets and composition functions) in $U$ for uniformly small values of $\epsilon$.
@@
The language of algebraic circuits provides quantitative metrics to formalize generalization in terms of algorithmic complexity, since circuit structure directly corresponds to algorithmic properties. In particular, generalization of AI systems can be quantified in terms of circuit divergence – the divergence of circuit parameters between algebraic circuits employed during training and testing (Fig. 3). This emphasis on characterizing the algorithmic properties of problems complements prior work focused on establishing metrics of compositionality based on the degree of subgraph overlap in training and testing sets. One particularly relevant metric is maximum compound divergence (MCD) introduced by Keysers et al. 17, which measures the divergence of the frequency distribution of compositional subgraphs between the training and testing sets. While Keysers et al. 17 demonstrated the empirical relevance of MCD on generalization performance, here we emphasize the importance of designing training and testing partitions according to the distribution of algorithmic complexity (i.e., circuit) properties. This contrasts with MCD, which focused on providing a single summary statistic that, though empirically useful, does not distinguish between algorithmic properties. Specifically, designing benchmarks by manipulating circuit size (vs. depth) enables the characterization of generalization over algorithmic size (vs. time) complexity. This in turn can help practitioners isolate how specific architectural components of models map onto algorithmic capabilities: for example, circuit depth of a problem typically relates to the number of layers of a transformer, while circuit size relates to the degree of parallelization, such as context length size for transformers 48.)
Given a model $M$, we seek to characterize its generalization performance $M(C_{test}|\mathcal{C})$, where $C_{test}$ is a test circuit, and $\mathcal{C}=\{C_{1},...,C_{k}\}$ is a family of training circuits. We define circuit divergence as the difference between quantifiable circuit properties between train and test distributions. Here, we emphasize five important circuit properties to measure divergences: size $s$, depth $d$, the polynomial degree, the sampling of a field $\mathbb{F}$, and the number of variables $\lvert X\rvert$. By quantifying the properties of circuits in both the training and testing sets, circuit divergence can be explicitly measured along these dimensions. In the following sections, we provide several examples where tests of generalization can be constructed through the manipulation of circuit divergence.
Figure 3: Algorithmic capabilities of modern AI systems and architectures can be studied with algebraic circuits. A) A set of problems can be identified (and sampled) from algebraic circuits. B) Given a family of circuits, we can identify circuit divergences – the divergence of different circuit properties, like size and depth – to design train and test datasets to evaluate the algorithmic generalization capabilities of a model. C,D) We can evaluate a model (or set of models) across these circuit splits, and quantify their performance according to circuit divergences. This AI evaluation lifecycle allows us to iterate and refine hypotheses regarding the degree to which a model can generalize to a class of algorithms. (Note that circuit divergences can be measured beyond depth and size.)
@@
In this section, we illustrate the flexibility of algebraic circuits in designing meaningful AI benchmarks. To demonstrate the generality of algebraic circuits, we begin by providing benchmarks that are analogous to popular tests of compositional generalization: systematicity and productivity 9. We subsequently introduce novel, more challenging problems that can be used to evaluate more general types of algorithmic computation and abstraction. Finally, we demonstrate an approach to reformulate tests of algorithmic generalization for pretrained LLMs.
@@
There has been recent interest in using compositional paradigms to systematically study the generalization capabilities of machine learning models 9, 8, 49, 50, 24, 51, 52, 12, 53, 54, 46. Here we demonstrate direct links to common forms of compositional generalization using algebraic circuits.